### v1

# A Note on Inequalities for Three Domination Parameters

Abstract
In this short paper, we establish relations between the domination number $\gamma$, the total domination number $\gamma_{\_}t$, and the connected domination number $\gamma_{\_}c$ of a graph. In particular, we prove upper and lower bounds for $\gamma_{\_}t$ in terms of $\gamma$ and $\gamma_{\_}c$. Moreover, we propose the following conjecture: for every connected isolated-free graph $G$,
$\gamma_{\_}t(G)\geqslant\left\lceil\frac{3\gamma(G)+2\gamma_{\_}c(G)}{6}\right\rceil.$
As evidence to support the conjecture, we prove that the conjecture holds when $\gamma_{\_}t(G)=\gamma_{\_}c(G)$ and also, when $\gamma_{\_}t(G)=\gamma_{\_}c(G)-1$.
Keywords: domination, total domination, connected domination.
2020 Mathematics Subject Classification: 05C69.
1 Introduction
In this paper, we deal with finite undirected and simple graphs. For a graph $G$, let $V(G)$ and $E(G)$ respectively denote the vertex set and the edge set of $G$.
Let $S$ be a subset of the vertex set of graph $G$. $S$ is called a dominating set of $G$ if every vertex not in $S$ is adjacent to at least one vertex in $S$. The domination number of $F$, denoted by $\gamma(G)$, is the minimum cardinality of a dominating set of $F$. A total dominating set of $G$ is a set $S$ such that every vertex is adjacent to some vertex in $S$. Note that by the definition of a total dominating set, it is evident that a graph $G$ admits a total dominating set if and only if $G$ has no isolated vertices. The total domination number of $G$, denoted by $\gamma_{\_}t(G)$, is the minimum cardinality of a total dominating set of $G$. A connected dominating set $S$ of $G$ is a set such that every vertex not in $S$ is adjacent with at least one vertex in $S$ and the subgraph induced by $S$ is connected. The connected domination number $\gamma_{\_}c(G)$ is the minimum cardinality of a connected dominating set of $G$. Note that a graph admits a connected dominating set if and only if the graph is connected.
The theory of domination in graphs is one of the main research areas in graph theory. In 1979, Bollobás and Cockayne [1] observed that the total domination number of an isolate-free graph is squeezed between the domination number and twice the domination number.
Theorem 1 ([1]).
If $G$ is an isolated-free graph, then $\gamma(G)\leqslant\gamma_{\_}t(G)\leqslant 2\gamma(G)$.
It is obvious that a connected dominating set is a dominating set. Sampathkumar and Walikar [10] proved the following result.
Theorem 2 ([10]).
If $G$ is a connected graph, then $\gamma(G)\leqslant\gamma_{\_}c(G)\leqslant 3\gamma(G)-2$.
The following easy fact was observed in [7].
Fact 1.
If $G$ is a connected isolated-free graph and $\gamma(G)>1$, then $\gamma(G)\leqslant\gamma_{\_}t(G)\leqslant\gamma_{\_}c(G)$.
A natural question is: what relationships exist between $\gamma,\gamma_{\_}t$ and $\gamma_{\_}c$?
To our knowledge, little or nothing is known about the relations between $\gamma,\gamma_{\_}t$ and $\gamma_{\_}c$, (and moreover, any three other domination parameters in general). This motivates us to study the relationships between $\gamma(G),\gamma_{\_}t(G)$ and $\gamma_{\_}c(G)$ when $G$ is a connected isolated-free graph.
For an excellent treatment of fundamentals of domination in graphs and recent topics on bounds for these three domination parameters, we refer the reader to the following books [5, 4, 3].
2 Main results
In this section, we demonstrate some relationships and show that these results are best possible.
Theorem 3.
If $G$ is a connected isolated-free graph and $\gamma(G)>1$, then
$\gamma_{\_}t(G)\leqslant\gamma(G)+\frac{1}{2}\gamma_{\_}c(G).$
Moreover, this bound is tight.
Proof.
Let $G$ be a connected isolated-free graph with $\gamma(G)>1$. From Theorem 1 and Fact 1, we have $\gamma_{\_}t(G)\leqslant 2\gamma(G)$ and $\gamma_{\_}t(G)\leqslant\gamma_{\_}c(G)$, respectively. By adding these two inequalities, we get $2\gamma_{\_}t(G)\leqslant 2\gamma(G)+\gamma_{\_}c(G)$ and the result follows immediately. In Figure 1, it can easily be checked that $\gamma(H)=5$ and $\gamma_{\_}t(H)=\gamma_{\_}c(H)=10$, which proves that the bound is tight. ∎
Figure 1: Graph $H$.
Similarly, we can deduce the following result from Theorems 1 and 2.
Theorem 4.
If $G$ is a connected isolated-free graph, then
$\gamma_{\_}t(G)\leqslant 5\gamma(G)-\gamma_{\_}c(G)-2.$
Moreover, this bound is tight.
Equality in Theorem 4 is achieved for complete graphs. In fact, the following generalisation of Theorem 3 holds for all connected isolated-free graphs.
Theorem 5.
For any connected isolated-free graph $G$,
$\gamma_{\_}t(G)\leqslant\left\lceil\frac{2(\gamma(G)+\gamma_{\_}c(G))}{3}\right\rceil.$
Moreover, this bound is tight.
Proof.
Let $G$ be a connected isolated-free graph. If $\gamma(G)>1$, then it follows from Fact 1 that $2\gamma_{\_}t(G)\leqslant 2\gamma_{\_}c(G)$. So, by Theorem 1, we have
$2\gamma_{\_}t(G)+\gamma_{\_}t(G)\leqslant 2\gamma(G)+2\gamma_{\_}c(G)\Rightarrow\gamma_{\_}t(G)\leqslant\frac{2(\gamma(G)+\gamma_{\_}c(G))}{3}.$
On the other hand, if $\gamma(G)=1$, then $G$ has a dominant vertex. That is, $G$ has a vertex that dominates every other vertex of $G$. In that case, $\gamma_{\_}t(G)=2$. So, $\lceil\frac{2\times 2}{3}\rceil=2$. Furthermore, this proves that the bound is tight. ∎
Now we turn our attention to the lower bounds. First, we point out the following lower bound, which is a consequence of Theorems 1 and 2.
Proposition 6.
If $G$ is a connected isolated-free graph, then
$\gamma_{\_}t(G)\geqslant 2\gamma(G)-\gamma_{\_}c(G).$
Moreover, this bound is tight.
Figure 2: Graph $G^{\prime}$.
Remark 1.
Equality is achieved for the graph $G^{\prime}$ in Figure 2. Moreover, the bound in Proposition 6 is not strong, as can be observed from the upper bound in Theorem 2. We believe that a stronger lower bound has the same form as Theorem 5. So, we propose the following conjecture.
Conjecture 7.
For every connected isolated-free graph $G$,
$\gamma_{\_}t(G)\geqslant\left\lceil\frac{3\gamma(G)+2\gamma_{\_}c(G)}{6}\right\rceil.$ (1)
Remark 2.
Conjecture 7 holds trivially for a connected isolated-free graph $G$ with $\gamma(G)=1$. Moreover, there are examples of graphs for which equality holds in (1). So, Conjecture 7, if true, is best possible.
As evidence of Conjecture 7, we prove the following results.
Theorem 8.
Let $G$ be a connected isolated-free graph. Then Conjecture 7 holds if
  1. (a)
    $\gamma_{\_}t(G)=\gamma_{\_}c(G)$.
  2. (b)
    $\gamma_{\_}t(G)=\gamma_{\_}c(G)-1$.
Proof.
For assertion (a), we suppose to the contrary that $\gamma_{\_}t(G)<\left\lceil\frac{3\gamma(G)+2\gamma_{\_}c(G)}{6}\right\rceil$. Then $\gamma_{\_}t(G)<\frac{3\gamma(G)+2\gamma_{\_}c(G)}{6}$ because $\gamma_{\_}t(G)\in\mathbb{N}$. So, we get
$6\gamma_{\_}t(G)<3\gamma(G)+2\gamma_{\_}c(G)\Rightarrow 4\gamma_{\_}t(G)<3\gamma(G).$
Theorem 1 implies that $4\gamma(G)\leqslant 4\gamma_{\_}t(G)<3\gamma(G)$, which is a contradiction.
For assertion (b), the proof is similar to that of assertion (a), so we omit it. ∎
Lemma 9.
Let $G$ be a connected isolated-free graph such that $\gamma(G)=2$. Then $\gamma_{\_}t(G)=\gamma_{\_}c(G)$.
Proof.
Let $u$ and $v$ be vertices in a dominating set of $G$. If $u$ and $v$ are adjacent, then $\gamma(G)=\gamma_{\_}t(G)=\gamma_{\_}c(G)=2$. On the other hand, suppose that $u$ and $v$ are not adjacent. Then $u$ and $v$ are connected by a path on at most four vertices (this includes $u$ and $v$). Otherwise, there is a vertex, say $w$, that is not dominated by either $u$ or $v$. If the path has three vertices, then $\gamma_{\_}t(G)=\gamma_{\_}c(G)=3$. Otherwise, $\gamma_{\_}t(G)=\gamma_{\_}c(G)=4$. ∎
By combining Theorem 8 (a) and Lemma 9, we get the following result.
Proposition 10.
Conjecture 7 holds for connected isolated-free graph $G$ with $\gamma(G)=2$.
Now we prove that Conjecture 7 holds for cycle graphs, but before that we will need the following result.
Proposition 11 ([6]).
For $n>2,\gamma_{\_}t(C_{\_}n)=\begin{cases}\frac{n}{2}&\mbox{if}\;n\equiv 0(\mbox{mod}\;4),\\ \frac{n+2}{2}&\mbox{if}\;n\equiv 2(\mbox{mod}\;4),\\ \frac{n+1}{2}&\mbox{otherwise}.\end{cases}$
Proposition 12.
Conjecture 7 holds for cycles $C_{\_}n$.
Proof.
Let $G$ be a cycle on $n$ vertices. Then $\gamma(G)=\lceil\frac{n}{3}\rceil$ and $\gamma_{\_}c(G)=n-2$ [6]. If $n\equiv 0$ (mod $3$), then
$\left\lceil\frac{3\gamma(G)+2\gamma_{\_}c(G)}{6}\right\rceil=\left\lceil\frac{n+2(n-2)}{6}\right\rceil=\left\lceil\frac{n}{2}-\frac{2}{3}\right\rceil\leqslant\frac{n}{2}\leqslant\gamma_{\_}t(G).$
On the other hand, if $n\not\equiv 0$ (mod $3$), then
$\left\lceil\frac{3\gamma(G)+2\gamma_{\_}c(G)}{6}\right\rceil\leqslant\left\lceil\frac{3(\frac{n}{3}+1)+2(n-2)}{6}\right\rceil=\left\lceil\frac{3n-1}{6}\right\rceil.$
If $4$ divides $n$, then $n=4k$, for some $k\in\mathbb{N}$. Then we have
$\left\lceil\frac{3n-1}{6}\right\rceil=\left\lceil\frac{3\times 4k-1}{6}\right\rceil=\left\lceil 2k-\frac{1}{6}\right\rceil\leqslant 2k=\gamma_{\_}t(G)$
If $4$ does not divide $n$, then $n=4k+c$, for some $k\in\mathbb{N}$ and $c\in\{1,2,3\}$. Thus,
$\left\lceil\frac{3n-1}{6}\right\rceil=\left\lceil\frac{3(4k+c)-1}{6}\right\rceil=\left\lceil 2k+\frac{3c-1}{6}\right\rceil\leqslant\frac{4k+c+1}{2}=2k+\frac{c+1}{2}=\gamma_{\_}t(G).$
Recall that for graphs $G$ and $H$, the Cartesian product $G\Box H$ is the graph with vertex set $V(G)\times V(H)$ where two vertices $(u_{\_}1,v_{\_}1)$ and $(u_{\_}2,v_{\_}2)$ are adjacent if and only if either $u_{\_}1=u_{\_}2$ and $v_{\_}1v_{\_}2\in E(H)$ or $v_{\_}1=v_{\_}2$ and $u_{\_}1u_{\_}2\in E(G)$.
Proposition 13 ([8]).
[…]
Proposition 14 ([2]).
[…]
Proposition 15 ([9]).
For any $n\geqslant 4$, $\gamma_{\_}c(P_{\_}4\Box P_{\_}n)=2n-\lfloor\frac{n}{3}\rfloor$.
Proposition 16.
Conjecture 7 holds for
  1. (a)
    $P_{\_}3\Box P_{\_}n$ for $n\geqslant 3$.
  2. (b)
    $P_{\_}4\Box P_{\_}n$ for $n\geqslant 4$
Proof.
For assertion (a), $\gamma_{\_}t(P_{\_}3\Box P_{\_}n)=\gamma_{\_}c(P_{\_}3\Box P_{\_}n)=n$ [2, 9] and the conclusion follows from Theorem 8 (a). For assertion (b), we will first show that for $n\geqslant 7$
$\left\lfloor\frac{6n+8}{5}\right\rfloor\geqslant\left\lceil\frac{3(n+1)+2(2n-\lfloor\frac{n}{3}\rfloor)}{6}\right\rceil.$
Suppose not. Then we have
$\frac{6n+8}{5}-1\leqslant\left\lfloor\frac{6n+8}{5}\right\rfloor<\left\lceil\frac{3(n+1)+2(2n-\lfloor\frac{n}{3}\rfloor)}{6}\right\rceil\leqslant\frac{7n+3-2\lfloor\frac{n}{3}\rfloor}{6}+1,$
which implies that $\frac{1}{30}n<\frac{9}{10}-\frac{1}{3}\lfloor\frac{n}{3}\rfloor\leqslant\frac{7}{30}$ for $n\geqslant 7$, a contradiction. The case $n=4,5$ or $6$ is an easy verification, and we leave that to the reader. ∎
Finally, we end this section by proving the following weakening of Conjecture 7.
Theorem 17.
For every connected isolated-free graph $G$,
$\gamma_{\_}t(G)\geqslant\left\lceil\frac{3\gamma(G)+\gamma_{\_}c(G)}{6}\right\rceil.$
Proof.
Suppose not. Then $\gamma_{\_}t(G)<\left\lceil\frac{3\gamma(G)+\gamma_{\_}c(G)}{6}\right\rceil$, which implies that $\gamma_{\_}t(G)<\frac{3\gamma(G)+\gamma_{\_}c(G)}{6}$ because $\gamma_{\_}t(G)\in\mathbb{N}$. Applying Theorem 2, we get
$26\gamma_{\_}t(G)<3\gamma(G)+\gamma_{\_}c(G)\leqslant 3\gamma(G)+3\gamma(G)-2=6\gamma(G)-$ 2 .
Once again, Theorem 1 implies that $26\gamma(G)\leqslant 6\gamma_{\_}t(G)<6\gamma(G)-$ 2; a contradiction. ∎
3 Conclusion
In this paper, we established relationships between $\gamma(G),\gamma_{\_}t(G)$ and $\gamma_{\_}c(G)$ for a connected isolated-free graph $G$. We hope that another major direction of research on domination in graphs is suggested by the results and the conjecture presented in this paper.